OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7

OpenStudy (anonymous):

parallel lines have the same slope in this case the slope for y = 4x + 7 is 4 so m=4 (x1,y1)=(-2, 2) so use point slope form y-y1=m(x-x1)

OpenStudy (anonymous):

so its y = -4x - 6?

OpenStudy (anonymous):

@RyanL.

OpenStudy (anonymous):

No it's not/ first just plug in the values you know first ....

OpenStudy (anonymous):

let's do this one step at a time....first plug the values in....

OpenStudy (anonymous):

y=1 and x=2 (im sorry i just dont get it

OpenStudy (anonymous):

where did you get y=1 and x=2 from?

OpenStudy (anonymous):

sorry mistype, cuse (-2, 2) that meaans y=-2 and +=2

OpenStudy (anonymous):

Ok we have the equation in point slope form \[y-y_1=m(x-x_1)\] \[m=6\] we don't know what x and y are and we will never know. so that leaves us with \[x_1 \]and\[y_1 \] we know our point \[(-2,2)\] which is the same as \[(x_1,y_1)\] that means that\[x_1=-2 \\ y_1=2\]

OpenStudy (zpupster):

m=4

OpenStudy (anonymous):

So now we have all the information so we do is plug in first.\[y-y_1=m(x-x_1) \\\]\[m=4\\x_1=-2\\y_1=2\]

OpenStudy (anonymous):

m=4 is first

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I fixed it for the last post I made sorry about that.

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

\[y-2=4(x+2)\] all you have to do now is simplify this and you are done.

OpenStudy (anonymous):

THANK YOUUUUU!!!!!!!!!!!!!

OpenStudy (anonymous):

how do you simplify that tho? @RyanL.

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

ok let my do that, thanks again @RyanL.

OpenStudy (zpupster):

if you are going ot simplify__ i am going to give a little shortcut you do not have to worry about y- y1 or x1 -x just remember- parallel lines have same slope so y = 4x + b jsut plug in your x and y given 2=4(-2) + b solve for b and then plug that back in y=4x+b(the b you solved for)

OpenStudy (anonymous):

ok ill do that thanks @zpupster